OpenStudy (anonymous):

What is the vertex of the parabola given by the equation y = 9(x + 4)2 – 3?

OpenStudy (anonymous):

First distribute the 9. y = 9(x+4)2-3 y = 9x + 36(2-3) Now distribute the 36 y = 9x + 72 - 108

OpenStudy (anonymous):

okay thank you but how do i find the vertex from there? its asking for a point..

OpenStudy (anonymous):

I think what you're trying to write is y = 9(x+4)^2 - 3. The carrot before the '2' show's that the (x+4) is raised by the power of 2. The definition of a vertex is where the line crosses the axis. In this case, the lien crosses the y-axis at x=0. Plug that into the equation y(0) = 9(0+4)^2-3 = 9*16-3 = 144 -3 = 141

OpenStudy (anonymous):

Ah ok then I was wrong sorry!

OpenStudy (anonymous):

thats okayy thanks anyways:) that is what i mean to write with the ^2 but the answer choices are points and i dont know how to get them from the simplified equation?

OpenStudy (anonymous):

why don't you give me the multiple choice answers... we could just plug in the x and y points ;)

OpenStudy (anonymous):

ha im stupid i didnt even think of that but the choices are (–4, –3) (–3, –4) (–3, 4) (4, –3)

OpenStudy (anonymous):

lol its ok... so lets plug in the first coordinates x = -4 and y = -3

OpenStudy (anonymous):

that one works:) thanks for your help!!

OpenStudy (anonymous):

You're welcome!